What volume of oxygen is produced from the decomposition of 2 moles of KCLO3 at s.t.p?
1
Expert's answer
2020-07-07T14:36:52-0400

2KClO3 -> 2KCl + 3O2

n(O2) = 2n (KClO3)/3= 4/3 mol

V (O2) = nRT/P = 4/3*0.0821*300/0.763= 43L